A thermometer reading 77∘77∘ F is placed in an oven preheated to a constant temperature. Through a glass window in the oven door, an observer records that the thermometer reads 279∘279∘ F after 1212 minute and 366∘366∘ F after 1 minute. How hot is the oven?
Answer #1

Gren that Let The Temperature of orent Tombe temperature of medium Then by Newtons bew of warning is dI =*(T-Ton) » Te T = mt279-Tm 366 - Tom y e 3660 219-m Bono, ekk = 279-Tm 77- TM Bom 4G, 366-m 274-Tm 279-TO 47-Tom فقا (Sth-C-TM) =(279-T) 26182 -3